The length of a rectangle exceeds the width by 2cm. If the diagonal is 10cm long, find the width of the rectangle.
Found 2 solutions by jim_thompson5910, swincher4391:
Answer by jim_thompson5910(35256)

L^2 + W^2 = D^2


(W+2)^2 + W^2 = 10^2


W^2 + 4W + 4 + W^2 = 100


W^2 + 4W + 4 + W^2 - 100 = 0


2W^2 + 4W - 96 = 0


2(W^2 + 2W - 48) = 0


2(W + 8)(W - 6) = 0


W + 8 = 0 or W - 6 = 0


W = -8 or W = 6


Throw out the negative solution to get W = 6


So the width is 6 cm and the length is 6+2 = 8 cm


If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim

Answer by swincher4391(1107)

Let l and w be the length and width respectively. Let d be the diagonal.
l = w+2
Then look at this diagram:
Photobucket
Notice we have a right triangle.
Then w^2 + (w+2)^2 = (10)^2
w^2 + w^2 + 4w + 4 = 100
2w^2 + 4w -96 = 0
2(w^2 + 2w - 48) = 0
2(w+8)(w-6) = 0
w = -8 and w = 6
Can't have a negative width, so w=6.
Thus our width is 6cm.